OpenStudy (anonymous):

A solution of 78% alcohol is to be mixed with a solution of 23% alcohol to form 165 L of a 58% solution. How many liters of the 78% solution must be used?

OpenStudy (kropot72):

Amount of alcohol in 58% solution = 165 * 0.58 Let amount of 78% solution = x Then amount of 23% solution = 165 - x 0.78x + 0.23(165 - x) = 165 * 0.58 Now solve the last equation to find x. Can you do that?

OpenStudy (anonymous):

not much different from the solution i wrote, is it?

OpenStudy (anonymous):

what is the answer because i got 2 different answers

OpenStudy (kropot72):

@mikeneedshelp1 What are your two answers?